# Self-Rewarding Language Models

### v1

@@ Abstract
We posit that to achieve superhuman agents, future models require superhuman feedback in order to provide an adequate training signal. Current approaches commonly train reward models from human preferences, which may then be bottlenecked by human performance level, and secondly these separate frozen reward models cannot then learn to improve during LLM training. In this work, we study Self-Rewarding Language Models, where the language model itself is used via LLM-as-a-Judge prompting to provide its own rewards during training. We show that during Iterative DPO training that not only does instruction following ability improve, but also the ability to provide high-quality rewards to itself. Fine-tuning Llama 2 70B on three iterations of our approach yields a model that outperforms many existing systems on the AlpacaEval 2.0 leaderboard, including Claude 2, Gemini Pro, and GPT-4 0613. While only a preliminary study, this work opens the door to the possibility of models that can continually improve in both axes.
@@ 1 Introduction
We thus introduce Self-Rewarding Language Models, agents that both (i) act as instruction following models generating responses for given prompts; and (ii) can generate and evaluate new instruction following examples to add to their own training set. We train these models using an Iterative DPO framework similar to that recently introduced in Xu et al. (2023). Starting from a seed model, as shown in Figure 1, in each iteration there is a process of Self-Instruction creation whereby candidate responses are generated by the model for newly created prompts, and are then assigned rewards by that same model. The latter is implemented via LLM-as-a-Judge prompting, which can also be seen as an instruction following task. A preference dataset is built from the generated data, and the next iteration of the model is trained via DPO.
@@ 2 Self-Rewarding Language Models
Our overall self-alignment procedure is an iterative one, which proceeds by building a series of such models, with the aim that each improves over the last. Importantly, because the model can both improve its generation ability, and it acts as its own reward model through the same generation mechanism, this means the reward model itself can improve through these iterations, deviating from standard practices where the reward model is fixed (Ouyang et al., 2022). We believe this can increase the ceiling of the potential for self-improvement of these learning models going forward, removing a constraining bottleneck.
@@ 2.1 Initialization
Seed instruction following data:
@@
Seed LLM-as-a-Judge instruction following data:
We also assume we are provided a seed set of (evaluation instruction prompt, evaluation result response) examples which can also be used for training. While this is not strictly necessary, as the model using IFT data will already be capable of training an LLM-as-a-Judge, we show that such training data can give improved results. In this data, the input prompt asks the model to evaluate the quality of a given response to a particular instruction. The provided evaluation result response consists of chain-of-thought reasoning (a justification), followed by a final score (in our experiments out of 5). The format we choose for these prompts is given in Figure 2. This thus serves as training data for the LLM to perform the role of a reward model. Subsequently this will be referred to as Evaluation Fine-Tuning (EFT) data.
@@ 2.2 Self-Instruction Creation
    Generate a new prompt: We generate a new prompt $x_{i}$ using few-shot prompting, sampling prompts from the original seed IFT data, following the approach of Wang et al. (2022) and Honovich et al. (2023).
@@
    Evaluate candidate responses: Finally, we use the LLM-as-a-Judge ability of our same model to evaluate its own candidate responses with scores $r_{i}^{n}\in[0,5]$ (see Figure 2).
@@ 2.3 Instruction Following Training
After performing the self-instruction creation procedure, we can then augment the seed data with additional examples for training, which we refer to as AI Feedback Training (AIFT) data. We try two variants of such feedback:
  • Preference pairs: We construct training data of the form (instruction prompt $x_{i}$, winning response $y_{i}^{w}$, losing response $y_{i}^{l}$). To form the winning and losing pair we take the highest and lowest scoring responses from the $N$ evaluated candidate responses (see Section 2.2), following Xu et al. (2023), discarding the pair if their scores are the same. These pairs can be used for training with a preference tuning algorithm. We use DPO (Rafailov et al., 2023).
  • Positive examples only: In this variant, we add additional examples of (instruction prompt, response) curated by the model to the seed set for supervised fine-tuning, following other approaches (Li et al., 2023a; Adolphs et al., 2023; Gulcehre et al., 2023), rather than constructing preference data. In this setup we only add examples where the candidate response was evaluated to give a perfect score of $r_{i}^{n}=5$.
While we report the results of both approaches in our experiments, we find that learning from preference pairs gives superior performance, and hence recommend that approach.
@@ 2.4 Overall Self-Alignment Algorithm
We thus define the models, and the training data they use as follows:
@@
This iterative training resembles the procedure used in Pairwise Cringe Optimization and Iterative DPO introduced in Xu et al. (2023); however, an external fixed reward model was used in that work.
@@ 3.1.1 Seed Training Data
We use the human-authored examples provided in the Open Assistant dataset (Köpf et al., 2023) for instruction fine-tuning. Following Li et al. (2023a) we use 3200 examples, by sampling only first conversational turns in the English language that are high-quality, based on their human annotated rank (choosing only the highest rank 0). In our experiments, we compare to a model fine-tuned from the base model using only this data via supervised fine-tuning, and refer to it as our SFT baseline.
@@
The Open Assistant data also provides multiple ranked human responses per prompt from which we can construct evaluation fine-tuning data. We split this into train and evaluation sets, and use it to create LLM-as-a-Judge data. This is done by placing it in the input format given in Figure 2, which consists of the scoring criteria description, and the given instruction and response to be evaluated. [^1] Note, the prompt, derived from Li et al. (2023a), mentions “utilizing web search”, but our model is not actually capable of this action. For training targets, chain-of-thought justifications and final scores out of 5 are not directly provided, so we use the SFT baseline to generate such output evaluations for each input, and accept them into the training set if the ranking of their scores agrees with the human rankings in the dataset. This results in 1775 train and 531 evaluation examples (which do not overlap with the IFT data).
@@ 3.1.2 Evaluation Metrics
We evaluate head-to-head performance between various models using GPT-4 (Achiam et al., 2023) as an evaluator over 256 test prompts derived from various sources following Li et al. (2023a) using the AlpacaEval evaluation prompt (Li et al., 2023b). We try the prompt in both orders comparing pairwise, and if the GPT-4 evaluations disagree we count the result as a tie. We also report results in the AlpacaEval 2.0 leaderboard format which is evaluated over 805 prompts, and compute the win rate against the baseline GPT-4 Turbo model based on GPT-4 judgments.
@@ 3.1.3 Training Details
The training hyperparameters we use are as follows. For SFT we use learning rate $5.5e{-6}$ which linearly decays to $1.1e{-6}$, batch size $16$ and dropout $0.1$. We only calculate the loss on target tokens instead of the full sequence. For DPO we use learning rate $1e{-6}$ which linearly decays to $1e{-7}$, batch size $16$, dropout $0.1$, and a $\beta$ value of 0.1. We perform early stopping by saving a checkpoint every 200 steps and evaluating generations using Claude 2 (Anthropic, 2023) on 253 validation examples derived from various sources following Li et al. (2023a). This is evaluated pairwise against the previous step’s generations using the AlpacaEval evaluation prompt format (Li et al., 2023b).
@@
To generate new prompts we used a fixed model, Llama 2-Chat 70B with 8-shot prompting, whereas the other parts of the creation pipeline (generating the response, and evaluating it) use the model being trained. For candidate response generation we sample $N=4$ candidate responses with temperature $T=0.7$, $p=0.9$. When evaluating candidate responses, as there is variance to these scores, in our experiments we also use sampled decoding (with the same parameters) and generate these evaluations multiple (3) times and take the average. We added 3,964 such preference pairs to form the AIFT($M_{1}$) dataset used to train $M_{2}$ via DPO, and 6,942 pairs to form AIFT($M_{2}$) used to train $M_{3}$.
@@ 3.2.1 Instruction Following Ability
We see a further gain in Iteration 3 over Iteration 2, with 47.7% wins for $M_{3}$ compared to only 12.5% for $M_{2}$ in a head to head evaluation. Similarly, the win rate over the SFT Baseline for $M_{3}$ increases to 62.5% wins vs. 9.8%, i.e. winning more often than the $M_{2}$ model did. Overall, we see large gains from $M_{2}$ to $M_{3}$ through training using the preference data AIFT($M_{2}$) provided by the reward model from Iteration 2.
@@
We evaluate our models on the AlpacaEval 2.0 leaderboard format, with results given in Table 1. We observe the same findings as in the head-to-head evaluations, that training iterations yield improved win rates over GPT4-Turbo, from 9.94% in Iteration 1, to 15.38% in Iteration 2, to 20.44% in Iteration 3. Our Iteration 3 model outperforms many existing models in this metric, including Claude 2, Gemini Pro, and GPT4 0613. We show some selected models from the leaderboard in the table. We note that many of those competing models contain either proprietary alignment data (which is typically large, e.g., over 1M annotations in (Touvron et al., 2023)) or use targets that are distilled from stronger models. In contrast, our Self-Rewarding model starts from a small set of seed data from Open Assistant, and then generates targets and rewards from the model itself for further iterations of training.
Preference optimization outperforms augmenting with positive examples only
We also tried the alternative self-training procedure of adding high-quality self-instruction creation examples to supervised fine-tuning (without preference optimization). Unfortunately we could not find a configuration where this approach helped. For example, adding 11,254 such examples that scored 5 out of 5, and optimizing the mixing weight in training, still yielded a head to head with the SFT Baseline of 29% wins vs 30% wins, i.e., no improvement.
Data distribution analysis
We perform a t-SNE (Van der Maaten and Hinton, 2008) visualization of the IFT, EFT and AIFT($M_{1}$) data, shown in Section A.1. We find good overlap between the IFT and AIFT($M_{1}$) examples, which is desired, while the EFT examples lie in a different part of the embedding space. We observe that generations from $M_{1}$ have an average length of 1092, for $M_{2}$ they are 1552, and for $M_{3}$ they are 2552, so the model is learning to generate longer responses, which we note may be a factor in relative performance.

### v3

@@ Abstract
We posit that to achieve superhuman agents, future models require superhuman feedback in order to provide an adequate training signal. Current approaches commonly train reward models from human preferences, which may then be bottlenecked by human performance level, and secondly these separate frozen reward models cannot then learn to improve during LLM training. In this work, we study Self-Rewarding Language Models, where the language model itself is used via LLM-as-a-Judge prompting to provide its own rewards during training. We show that during Iterative DPO training that not only does instruction following ability improve, but also the ability to provide high-quality rewards to itself. Fine-tuning Llama 2 70B on three iterations of our approach yields a model that outperforms many existing systems on the AlpacaEval 2.0 leaderboard, including Claude 2, Gemini Pro, and GPT-4 0613. While there is much left still to explore, this work opens the door to the possibility of models that can continually improve in both axes.
@@ 1 Introduction
We thus introduce Self-Rewarding Language Models, that both (i) act as instruction following models generating responses for given prompts; and (ii) can generate and evaluate new instruction following examples to add to their own training set. We train these models using an Iterative DPO framework similar to that recently introduced in Xu et al. (2023). Starting from a seed model, in each iteration there is a process of Self-Instruction creation whereby candidate responses are generated by the model for newly created prompts, and are then assigned rewards by that same model. The latter is implemented via LLM-as-a-Judge prompting, which can also be seen as an instruction following task. A preference dataset is built from the generated data, and the next iteration of the model is trained via DPO, see Figure 1.
@@ 2 Self-Rewarding Language Models
Our overall self-alignment procedure is an iterative one, which proceeds by building a series of such models, with the aim that each improves over the last. Importantly, because the model can both improve its generation ability, and act as its own reward model through the same generation mechanism, this means the reward model itself can improve through these iterations, deviating from standard practices where the reward model is fixed (Ouyang et al., 2022). We believe this can increase the ceiling of the potential for self-improvement of these learning models going forward, removing a constraining bottleneck.
@@ 2.1 Initialization
Seed instruction following data
@@
Seed LLM-as-a-Judge instruction following data
We also assume we are provided a seed set of (evaluation instruction prompt, evaluation result response) examples which can also be used for training. While this is not strictly necessary, as the model using IFT data will already be capable of training an LLM-as-a-Judge, we show that such training data can give improved performance (see Appendix A.3 for supporting results). In this data, the input prompt asks the model to evaluate the quality of a given response to a particular instruction. The provided evaluation result response consists of chain-of-thought reasoning (a justification), followed by a final score (in our experiments out of 5). The exact prompt format we chose is given in Figure 2, which instructs the LLM to evaluate the response using five additive criteria (relevance, coverage, usefulness, clarity and expertise), covering various aspects of quality. Subsequently this will be referred to as Evaluation Fine-Tuning (EFT) data.
@@ 2.2 Self-Instruction Creation
    Generate a new prompt: We generate a new prompt $x_{i}$ using few-shot prompting, sampling prompts from the original seed IFT data, following the approach of Wang et al. (2023) and Honovich et al. (2023). [^1] In our main experiments, responses and rewards, items (2) and (3), are generated by the model we have trained, but generating prompts is actually done by a model fixed in advance. However, we show that prompts can also be generated by the newly trained model in each iteration in Appendix A.5.
@@
    Evaluate candidate responses: Finally, we use the LLM-as-a-Judge ability of our same model to evaluate its own candidate responses with scores $r_{i}^{n}\in[0,5]$ (exact prompt given in Figure 2).
@@ 2.3 Instruction Following Training
After performing the self-instruction creation procedure, we can augment the seed data with additional examples for training, which we refer to as AI Feedback Training (AIFT) data.
To do this, we construct preference pairs, which are training data of the form (instruction prompt $x_{i}$, winning response $y_{i}^{w}$, losing response $y_{i}^{l}$). To form the winning and losing pair we take the highest and lowest scoring responses from the $N$ evaluated candidate responses (see Section 2.2), following Xu et al. (2023), discarding the pair if their scores are the same. These pairs can be used for training with a preference tuning algorithm. We use DPO (Rafailov et al., 2023).
@@ 2.4 Overall Self-Alignment Algorithm
We define the models, and the training data they use as follows:
@@
This iterative training resembles the procedure used in Pairwise Cringe Optimization and specifically is termed Iterative DPO, introduced in Xu et al. (2023); however, an external fixed reward model was used in that work.
@@ 3.1.1 Seed Training Data
We use the human-authored examples provided in the Open Assistant dataset (Köpf et al., 2023) for instruction fine-tuning. Following Li et al. (2024) we use 3200 examples, by sampling only first conversational turns in the English language that are high-quality, based on their human annotated rank (choosing only the highest rank 0). In our experiments, we compare to a model fine-tuned from the base model using only this data via supervised fine-tuning, and refer to it as our SFT baseline.
@@
The Open Assistant data also provides multiple ranked human responses per prompt from which we can construct evaluation fine-tuning data. We split this into train and evaluation sets, and use it to create LLM-as-a-Judge data. This is done by placing it in the input format given in Figure 2, which consists of the scoring criteria description, and the given instruction and response to be evaluated. [^2] Note, the prompt, derived from Li et al. (2024), mentions “utilizing web search”, but our model is not actually capable of this action. For training targets, chain-of-thought justifications and final scores out of 5 are not directly provided, so we use the SFT baseline to generate such output evaluations for each input, and accept them into the training set if the ranking of their scores agrees with the human rankings in the dataset. We resample the training set by discarding some of the data that receives the most common score so that the scores are not too skewed, as we observe many samples receive a score of 4. This results in 1,630 train and 541 evaluation examples (which do not overlap with the IFT data).
@@ 3.1.2 Evaluation Metrics
We evaluate head-to-head performance between various models using GPT-4 (Achiam et al., 2023) as an evaluator over 256 test prompts (which we refer to as IFT test data) derived from various sources following Li et al. (2024) using the AlpacaEval evaluation prompt (Li et al., 2023). We try the prompt in both orders comparing pairwise, and if the GPT-4 evaluations disagree we count the result as a tie. We also perform a similar evaluation with humans (authors). We additionally report results in the AlpacaEval 2.0 leaderboard format which is evaluated over 805 prompts, and compute the win rate against the baseline GPT-4 Turbo model based on GPT-4 judgments. Further, we report results on MT-Bench (Zheng et al., 2023b) a set of challenging multi-turn questions in various categories from math and coding to roleplay and writing, which uses GPT-4 to grade the model responses out of 10. Finally we also test the models on a set of 9 NLP benchmarks: ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), SIQA (Sap et al., 2019), PIQA (Bisk et al., 2020), GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021), OBQA (Mihaylov et al., 2018) and NQ (Kwiatkowski et al., 2019).
@@ 3.1.3 Training Details
The training hyperparameters we use are as follows. For SFT we use learning rate $5.5e{-6}$ which decays (cosine) to $1.1e{-6}$ at the end of training, batch size $16$ and dropout $0.1$. We only calculate the loss on target tokens instead of the full sequence. For DPO we use learning rate $1e{-6}$ which decays to $1e{-7}$, batch size $16$, dropout $0.1$, and a $\beta$ value of 0.1. We perform early stopping by saving a checkpoint every 200 steps and evaluating generations using Claude 2 (Anthropic, 2023) on 253 validation examples derived from various sources following Li et al. (2024). This is evaluated pairwise against the previous step’s generations using the AlpacaEval evaluation prompt format (Li et al., 2023).
@@
To generate new prompts we use a fixed model, Llama 2-Chat 70B with 8-shot prompting following Self-Instruct (Wang et al., 2023), where we sample six demonstrations from the IFT data and two from the model generated data, and use decoding parameters T = 0.6, p = 0.9. We use their prompt template for non-classification tasks and apply the same filtering techniques, including the ROUGE-L (Lin, 2004) similarity check, keyword filtering, and length filtering. Except for the prompt generation part, the other parts of the creation pipeline (generating the response, and evaluating it) use the Self-Rewarding model being trained. For candidate response generation we sample $N=4$ candidate responses with temperature $T=0.7$, $p=0.9$. When evaluating candidate responses, as there is variance to these scores, in our experiments we also use sampled decoding (with the same parameters) and generate these evaluations multiple (3) times and take the average. We added 3,964 such preference pairs to form the AIFT($M_{1}$) dataset used to train $M_{2}$ via DPO, and 6,942 pairs to form AIFT($M_{2}$) used to train $M_{3}$.
@@ 3.2.1 Instruction Following Ability
We see a further gain in Iteration 3 over Iteration 2, with 47.7% wins for $M_{3}$ compared to only 12.5% for $M_{2}$ in a head to head evaluation. Similarly, the win rate over the SFT Baseline for $M_{3}$ increases to 62.5% wins vs. 9.8%, i.e., winning more often than the $M_{2}$ model did. Overall, we see large gains from $M_{2}$ to $M_{3}$ through training using the preference data AIFT($M_{2}$) provided by the reward model from Iteration 2.
@@
We evaluate our models on the AlpacaEval 2.0 leaderboard format, with results given in Table 1. We observe the same findings as in the head-to-head evaluations, that training iterations yield improved win rates, in this case over GPT4-Turbo, from 9.94% in Iteration 1, to 15.38% in Iteration 2, to 20.44% in Iteration 3. Our Iteration 3 model outperforms many existing models in this metric, including Claude 2, Gemini Pro, and GPT4 0613. We show some selected models from the leaderboard in the table. We note that many of those competing models contain either proprietary alignment data (which is typically large, e.g., over 1M annotations in Touvron et al. (2023)) or use targets that are distilled from stronger models. In contrast, our Self-Rewarding model starts from a small set of seed data from Open Assistant, and then generates targets and rewards from the model itself for further iterations of training.
@@
Figure 4: AlpacaEval win rate breakdown for instruction categories (full names given in Appendix). Self-Rewarding models give gains across several topics, but tend to e.g. give less gains on mathematics and reasoning tasks.
Fine-grained analysis
As described earlier, the overall performance of the model in AlpacaEval improves with each iteration of training. It would be interesting to break down the overall performance improvement to see exactly what type of tasks these improvements come from. Therefore, we cluster the instructions in AlpacaEval test set into different groups based on three perspectives: (1) instruction category (2) instruction complexity (3) expected response length. We achieve this by using GPT-4. The detailed statistical information of the breakdown and the prompting techniques we used for getting this breakdown can be found in Appendix A.6. Results for the instruction category are given in Figure 4, and the other two in Appendix Figure 11. From the results we can conclude that (i) Self-Rewarding models can substantially improve the win rate in most categories, but there are some tasks for which this approach does not improve, such as mathematics and logical reasoning, indicating that our current training approach mainly allows the models to better utilize their existing knowledge. (ii) Through Self-Rewarding model training, the model’s win rate increases on almost all tasks of different complexity, and especially on slightly more difficult tasks (complexity of 5, 6, 7 out of 10). (iii) The models also show a steady increase in the win rate on tasks with instructions with different expected response lengths.
Data distribution analysis
We perform a t-SNE (Van der Maaten and Hinton, 2008) visualization of the IFT, EFT and AIFT($M_{1}$) data, shown in Appendix A.1. We find good overlap between the IFT and AIFT($M_{1}$) examples, which is desired, while the EFT examples lie in a different part of the embedding space, which can help explain why they would not affect IFT performance. We observe that generations from $M_{1}$ on AlpacaEval have an average length of 1092, for $M_{2}$ they are 1552, and for $M_{3}$ they are 2552, so the model is learning to generate longer responses, which we note may be a factor in relative performance.
Figure 5: Human evaluation results. Iterations of Self-Rewarding ($M_{1}$, $M_{2}$ and $M_{3}$) provide progressively better head-to-head win rates compared to the SFT baseline, in agreement with the automatic evaluation results.
Table 2: MT-Bench Results (on a scale of 10). Self-Rewarding iterations yield improving scores across various categories. Math, code & reasoning performance and iteration gains are smaller than for other categories, likely due to the makeup of the Open Assistant seed data we use.
Overall Math, Code Humanities, Extraction,
Score & Reasoning STEM, Roleplay & Writing
SFT Baseline 6.85 3.93 8.60
$M_{1}$ 6.78 3.83 8.55
$M_{2}$ 7.01 4.05 8.79
$M_{3}$ 7.25 4.17 9.10
Table 3: NLP Benchmarks. Self-Rewarding models mostly tend to maintain performance compared to the Llama 2 70B base model and the SFT Baseline, despite being fine-tuned on very different instruction-following prompts.
ARC ()($\uparrow$)( ↑ )
challenge
HellaSwag
()($\uparrow$)( ↑ )
GSM8K
()($\uparrow$)( ↑ )
MMLU
()($\uparrow$)( ↑ )
NQ
()($\uparrow$)( ↑ )
Llama 2 57.40 85.30 56.80 68.90 25.30
SFT Baseline 55.97 85.17 50.72 69.76 34.35
$M_{1}$ 57.51 84.99 60.27 69.34 35.48
$M_{2}$ 54.51 84.27 59.29 69.31 33.07
$M_{3}$ 53.13 83.29 57.70 69.37 31.86
Human evaluation
To examine whether human judgments align with automatic evaluation results, we conduct human evaluations that compare SFT baseline generations with the generations from each iteration of Self-Rewarding training, i.e., models $M_{1}$, $M_{2}$, and $M_{3}$. Specifically, we randomly select 50 instructions from the IFT test set. Each instruction corresponds to three pairs of generations (i.e., baseline vs. $M_{1}$, baseline vs. $M_{2}$, baseline vs. $M_{3}$). For each pair of generations, we assign them to three different annotators (blind evaluation performed by the authors) to make a pairwise judgment, and take a majority vote to decide which generation is better. The human evaluation results are shown in Figure 5. We find that Self-Rewarding models from later iterations show a larger advantage over the SFT baseline model, which is consistent with GPT-4’s judgments, and demonstrates the effectiveness of our iterative training procedure.
MT-Bench performance further validates these results
We report performance on MT-Bench in Table 2 for the SFT baseline and iterations of the Self-Rewarding model. We again see improvements across the iterations of training from $M_{1}$ to $M_{3}$, from 6.78 (out of 10) up to 7.25, with larger relative gains in the humanities, STEM, roleplay, writing and extraction categories, and smaller gains in the math, code and reasoning categories. We expect that the latter is due to the seed prompts we use from Open Assistant tending to underemphasize the reasoning-based tasks. We note also that these improvements are in spite of our method using and constructing prompts that only involve a single turn, given the MT-Bench benchmark itself is a multi-turn evaluation.
Self-rewarding models did not lose ability on NLP Benchmarks
As shown in Table 3, the performance of most NLP benchmark tasks evaluated are roughly similar to the baselines, with further detailed results on more datasets given in Appendix Table 9 that follow the same pattern. We hypothesize that given that our training data (seed data and synthetically generated data) are based on the Open Assistant prompts which may not be especially relevant to skills needed in the Table 3 tasks, it is expected that the task performance stays roughly similar, or may even drop. For example, in InstructGPT training (Ouyang et al., 2022) they found that “during RLHF fine-tuning, we observe performance regressions compared to GPT-3 on certain public NLP datasets” which they refer to as an “alignment tax.” A clear future direction is to extend the self-rewarding paradigm to these types of tasks, by relying not only on seed prompts from Open Assistant, but also on seed prompts found in a larger variety of datasets.
3.2.2 Reward Modeling Ability
Reward modeling evaluation results are provided in Table 4.
EFT augmentation improves over SFT baseline
Firstly, we find that adding Evaluation Fine-Tuning (EFT) data into training, which gives examples to the model of how to act as an LLM-as-a-Judge, naturally improves its performance compared to training with Instruction Fine-Tuning (IFT) data alone. IFT data covers a wide range of general instruction tasks, and so does endow the SFT Baseline with the ability to evaluate responses; however, EFT data gives more examples of this specific task. We find improvements across all five metrics measured when using IFT+EFT vs. IFT alone, e.g., the pairwise accuracy agreement with humans increases from 65.1% to 78.7%.